OpenStudy (anonymous):

The area of a rectangle is x2 - 2x - 15 and the length of the rectangle is x + 3. Find the width of the rectangle. A. x – 5 B. x2 + 15 C. x + 5 D. (x – 2)(x + 3)

OpenStudy (anonymous):

i put b?

jimthompson5910 (jim_thompson5910):

Length * Width = Area (x+3) * (x^2 + 15) = x^2 - 2x - 15 is that last equation true for all values of x?

OpenStudy (tkhunny):

Can't be. The area is Second Degree. \(x^{2} - 2x - 15 = SOMETHING*(x+3)\) I'm guessing that left hand side can be factored.

OpenStudy (anonymous):

you can factor the equation and you would find the width of the rectangle

OpenStudy (anonymous):

and b its wrong

OpenStudy (anonymous):

i figured ...

OpenStudy (anonymous):

good job :)

OpenStudy (anonymous):

since i got it wrong... =*(

OpenStudy (anonymous):

c?

jimthompson5910 (jim_thompson5910):

what do you get when you expand out (x+3)(x+5)

OpenStudy (anonymous):

Its wrong

OpenStudy (anonymous):

x2 - 2x - 15=(x+3)(x_ _) on the blank you have to find the (+ or -) and a number

OpenStudy (anonymous):

@hinarauf80

OpenStudy (anonymous):

ummm -?

OpenStudy (anonymous):

a or d

OpenStudy (anonymous):

its a

OpenStudy (anonymous):

ok man thanks

OpenStudy (anonymous):

YW!